This section contains 5 questions. The answer to each of the questions is a single digit integer ranging from 0 to 9.


 


While measuring length of an object it was observed that the zero of the vernier lies between 1.4 and 1.5 of the main scale and the fifth vernier division coincides with a main scale division. If the length of the object measured is l, the value of (l – 1.4) in terms of the least count C of the instrument is yC. Find y.


 

Grade:upto college level

1 Answers

Radhika Batra
10 years ago

5

 

        Let x =  – 1.4

 

Then 1.4 + x + 5V = 1.4 + 5 S, where S denotes the length of one main scale division, and V denotes the length of one vernier scale division.

 

Hence, x = 5 S – 5 V = (1S – 1V) = 5C

 

\ y = 5